OpenStudy (anonymous):

Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (radicalsniper):

X = Xo + Vo*t +0.5(a)(t)^2 where X = the final distance, Xo is the initial distance, Vo = to the initial velocity, a and t is acceleration and time respectively. Since the question is how long does it take until the bikes are 45 miles apart, We will add the two velocities. X= Velocity of the first bike * time + Velocity of the second bike * time . Xo = 0 and acceleration = 0. 45 = 10mph(t) + 5mph(t) t= 3 hours.